OpenStudy (goformit100):

The radius of a circle is increasing at the rate of 0.7 cm/s. What is the rate of increase of its circumference?

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

\[\huge C = 2\pi r\]\[\huge \frac{dC}{dr}=2\pi\]

OpenStudy (anonymous):

\[\frac{ dr }{ dt }=0.7 \ cm/s\]\[C=2 \pi r\]\[\frac{ dC }{ dt }=2 \pi\]\[\bf \frac{ dC }{ dr }=\frac{ \frac{ dC }{ dt } }{ \frac{ dr }{ dt } }\] @goformit100

OpenStudy (anonymous):

Plug everything in again and solve.

terenzreignz (terenzreignz):

\[\large \frac{dC}{dt} =_{?} 2\pi \] @genius12 I dispute this...

OpenStudy (goformit100):

Thank you Sir

OpenStudy (anonymous):

Sorry @terenzreignz I forgot to include the dr/dt in dC/dt lol.

OpenStudy (anonymous):

@goformit100 \[\frac{ dC }{ dt }=2\pi \frac{ dr }{ dt }\] Please make this correction.

OpenStudy (goformit100):

Thanks @genius12